# Freezing-in the Axiverse

### v1

@@ 3.2 Couplings to Fermions
In this section, we consider an EFT where axions couple to fermions at tree-level, in addition to the gauge sector, motivated by the DFSZ scenario Dine:1981rt; Zhitnitsky:1980tq. Those couplings can be written as in (18), which we reproduce here for convenience,
@@ 3.2.1 Anarchy
3.2.2 Froggatt-Neilsen Texture
We next assume that the axion couplings to the fermion sector are such that the SM flavor puzzle is solved by the Froggatt-Neilsen (FN) mechanism Froggatt:1978nt at a scale $\Lambda_{\rm FN}\gg f_{a}$. In the FN setup, the SM is extended with a global $U(1)_{\cal H}$ horizontal symmetry which is spontaneously broken by the VEV $\langle S\rangle$ of a new complex scalar $S$. The SM Yukawas arise solely from powers of $\epsilon\equiv\langle S\rangle/\Lambda_{\rm FN}$.
This idea can be generalized by considering a texture Leurer:1992wg; Leurer:1993gy, the set of charges of the SM fermions under $U(1)_{\cal H}$. In this work, we consider such an example of a texture, but we take a bottom-up approach and do not specify the UV degrees of freedom that provide such charges. We adopt a convention where $S$ has charge 1 and the SM Higgs has charge 0 under $U(1)_{\cal H}$, we can express the SM Yukawa couplings
@@ 3.2.2 Froggatt-Neilsen Texture
where $\kappa$’s are some ${\cal O}(1)$ numbers. In most implementations, the ratio $\langle S\rangle/\Lambda_{\rm FN}$ is identified with $\epsilon\equiv\sin\theta_{c}\approx 0.23$, and a texture is identified that reproduces the SM Yukawas at low energies. To conserve charge, we see that $n_{nm}^{d}=|{\cal H}(Q_{n})-{\cal H}(d_{m})|$ and $n_{nm}^{u}=|{\cal H}(Q_{n})-{\cal H}(u_{m})|$. Note that if the argument of the absolute value is negative, then instead $S^{\dagger}$ is coupled to the SM fermions. Various choices of texture are possible, see, e.g.Leurer:1992wg; Leurer:1993gy. Once the FN charges are assigned to the SM fermions, they also dictate the sizes of the couplings of those fermions to other fields, including new physics scenarios like the axion couplings. In particular, the structure of the $\phi_{i}$ coupling to the SM fermions requires $c_{F,nm}\propto\epsilon^{|{\cal H}(\Psi_{F}^{n})-{\cal H}(\Psi_{F}^{m})|}$. Using the horizontal charge assignments given in Table 1 of Asadi:2023ucx, we find
@@
The numbers $X,Y$ take values $\pm 1$ and together with the other sign choices in $c_{u},c_{d},c_{e}$, there are $2^{5}$ possibilities for the FN charges. For an example benchmark, we choose X=Y=1X=-Y=-1, the positive sign in $c_{u},c_{d}$, and the negative sign in $c_{e}$; this choice is motivated by holomorphy of a SUSY UV completion of the axiverse EFT Asadi:2023ucx. This leads to
@@
Finally, we note that the angular mode of $S$ is a light pseudoscalar which could also freeze-in. In this work, we do not consider its contribution to $\Delta N_{\rm eff}$. In some models it is identified with the axion Ema:2016ops; Calibbi:2016hwq.
@@ 3.2.3 Minimal Flavor Violation
Our final assumption for the axion-fermion couplings does introduce correlations between the axions coupled to different SM fermions, so we will now find that there are less than 42 axion-fermion couplings at dimension five. We investigate in this context the principle of Minimal Flavor Violation (MFV) DAmbrosio:2002vsn, where the SM Yukawas transform as spurions under the $SU(3)^{5}$ flavor symmetry group. Then the axion-fermion couplings can be expanded in powers of the spurions, where $c_{F}$ is required to transform as $\mathbf{8}_{F}$ under $SU(3)_{\psi}\subset SU(3)^{5}$. It is useful to define the Hermitian matrices $X_{U}\equiv Y_{u}Y_{u}^{\dagger}\sim\mathbf{8}_{Q}$, $X_{D}\equiv Y_{d}Y_{d}^{\dagger}\sim\mathbf{8}_{Q}$, $U_{k}\equiv Y_{u}^{\dagger}X_{D}^{k-1}Y_{u}\sim\mathbf{8}_{u}$, $D_{k}\equiv Y_{d}^{\dagger}X_{U}^{k-1}Y_{d}\sim\mathbf{8}_{d}$, $L\equiv Y_{e}Y_{e}^{\dagger}\sim\mathbf{8}_{L}$, $E\equiv Y_{e}^{\dagger}Y_{e}\sim\mathbf{8}_{e}$. Using the Cayley-Hamilton theorem, we expand the couplings in only these matrices and remove related terms. We find
@@
where all expansion coefficients are expected to be $\mathcal{O}(1)$. In the limit of zero neutrino masses, the expansion in the lepton sector is complete. While we have not explicitly written the full expansion in the quark sector, it can be systematically computed Colangelo:2008qp. In our numerical computation, we use the full expansion without the truncation shown above. It is useful to work in a basis where one of the quark Yukawa couplings is diagonal. We label the up-basis the one where $Y_{u}={\rm diag}(y_{u},y_{c},y_{t})$ and $Y_{d}=V_{\rm CKM}{\rm diag}(y_{d},y_{s},y_{b})$. The down-basis is related by a $SU(3)_{Q}$ rotation by $V_{\rm CKM}^{\dagger}$. Explicitly, in either basis $c_{u}$, $c_{d}$, $c_{L}$, and $c_{e}$ are identical, while $c_{Q}^{\rm up}=V_{\rm CKM}c_{Q}^{\rm down}V_{\rm CKM}^{\dagger}$. We choose to work in the up-basis. We have explicitly checked that working in the down-basis leads to identical results.
@@ 4.1 Effects of Mass Mixing
For simplicity, consider a two-axion system, where the axions have masses $m_{1}$ and $m_{2}$. The mass basis does not coincide with the flavor basis generically, and therefore, the axion of one flavor can oscillate into another flavor. In analogy with neutrino oscillations in vacuum, the probability of a flavor oscillation in time $t$ is given by $\sin^{2}(\vartheta)\times\sin^{2}((m_{2}^{2}-m_{1}^{2})t/(4k))$ Raffelt:1996wa, where $\vartheta$ is the rotation angle between the mass basis and the flavor basis, and $k$ is the 3-momenta of both the axions. Here, we have treated both axions to be relativistic. Since we are interested in the production of these axions from the thermal bath, $k\sim T$, the temperature of the SM bath. Taking $t\sim 1/H$, at $T\gg{\rm GeV}$ and $m_{2},m_{1}\ll\rm{eV}$ (as we will justify shortly),
@@ 4.2 QCD Axion Cosmology
Depending on its mass, the QCD axion can either contribute as dark matter or dark radiation. Since our EFT treatment is valid for $T_{\rm RH}<f_{a}$, we only focus on DM production through the misalignment mechanism, which implies ParticleDataGroup:2024cfk
@@
for misalignment angle $\theta_{i}\lesssim 1$. The QCD axion can also be produced from the SM bath, either through pion scattering or through quarks and gluons, depending on the temperature. Ref. Notari:2022ffe places an upper bound $m_{\rm QCD}\leq 0.24$ eV at 95% CL, implying $f_{a}\geq 2.4\times 10^{7}$ GeV, by computing $a_{\rm QCD}$ production from pion scattering at temperatures $\lesssim 150$ MeV, the QCD crossover temperature. Above this temperature, but below the electroweak scale, perturbative computations are not fully trustworthy (see Notari:2022ffe and references therein). However, convergence of perturbative QCD improves at higher temperatures, and we expect the associated rates to be under better control for $T_{\rm RH}\gtrsim$ TeV, which is where we focus on in this work.
@@ 4.4.1 Production via Gauge Couplings
To estimate axion production from gauge couplings, we use the production rates based on the thermal field theory calculations in Salvio:2013iaa (for calculations at zero temperature, see Masso:2002np). These rates also determine the axion decoupling temperatures $T_{d}$, which is when the axion production rate equals the Hubble expansion rate. We evaluate $T_{d}$ for the three gauge interactions and show the result in Fig. 1, including the GUT scenario. We also use these thermal rates to obtain an approximate abundance of $\Delta N_{\rm eff}$,
@@ 4.4.2 Production via Fermionic Couplings
The associated freeze-in production can be computed analytically via solving the full Boltzmann equation for the axion phase space distribution function. [^3] For a general numerical treatment see, e.g.DEramo:2023nzt; DEramo:2024jhn. We can express the relevant processes for freeze-in production as,
@@
where we have defined a momentum-dependent ‘production rate’ $\Gamma(p,t)$. To solve this we note the physical momentum $p=q/a$ where $q$ is the comoving momentum. We then express $\mathfrak{f}_{a}(p,t)=\tilde{\mathfrak{f}}_{a}(q,a)$ and analogously for the other quantities. The Boltzmann equation then becomes a collection of ordinary differential equations for each comoving momentum DAgnolo:2017dbv,
@@ 4.4.3 Production via Dim-6 Higgs Coupling
However, the fact that there are two axions in the final state leads to a nonlinear integro-differential equation which must be solved numerically. We leave this to future work, and make the approximation that we can ignore the backwards process: $\mathfrak{f}_{a}(\boldsymbol{k}_{1},t)\mathfrak{f}_{a}(\boldsymbol{k}_{2},t)=0$. Clearly, this can only be justified in the regime that the axion does not thermalize; we ceiling $\Delta N_{\rm eff}$ at a thermal abundance. Following Edsjo:1997bg; Hall:2009bx; Blennow:2013jba, one can then solve the (integrated) Boltzmann equations for the axion number density,
@@
While this result matches parametrically with Ref. Bauer:2022rwf, its numerical value differs by several orders of magnitude. For a verification, we have checked that our computation matches with micrOMEGAs Alguero:2023zol numerical output.
@@ 5 Results for Benchmark EFTs
We also show the sensitivity of current and upcoming CMB experiments. The horizontal lines correspond to existing one-sided 95% upper limits $\Delta N_{\rm eff}^{95\%}$. SPT-Planck-ACT (SPA) corresponds to the constraint from the recent SPT-3G analysis SPT-3G:2025bzu when floating the helium abundance $Y_{P}$ in the fit, yielding $N_{\rm eff}=2.99_{-0.26}^{+0.22}$, so that $\Delta N_{\rm eff}^{95\%}\approx 0.31$. We note that without floating $Y_{P}$, we find $\Delta N_{\rm eff}^{95\%}\approx-0.04$ due to the low central values from Planck and ACT. For this reason, we adopt the more conservative result from floating $Y_{P}$. Indeed, the ACT results ACT:2025tim (including Planck lensing and BOSS BAO) yield $\Delta N_{\rm eff}^{95\%}\approx 0.04$. We show instead the result from restricting to physical parameter space with $\Delta N_{\rm eff}>0$: $\Delta N_{\rm eff}^{95\%}=0.17$, but in plots we shade only the region excluded by the SPA analysis. Finally, we show projections for Simons Observatory (SO), CMB-S4 (S4), and CMB-HD (HD) SimonsObservatory:2025wwn; CMB-S4:2016ple; MacInnis:2023vif.
@@
Figure 6: A summary of our results considering only dimension-5 couplings for $f_{a}=10^{12}$ GeV as a function of $T_{\rm RH}$ and different assumptions about the axion couplings. The existing and future CMB constraints are shown as horizontal lines: the combination of SPT-3G, Planck, and ACT (SPA, black solid), ACT alone (gray solid), SO (dashed), S4 (dot-dashed), and HD (dotted). The $\Delta N_{\rm eff}$ expected are also shown assuming anarchic couplings, a Froggatt-Neilsen texture, MFV, hadronic, and a hadronic GUT. The kink on the hadronic GUT (blue) line around $2\times 10^{11}$ GeV indicates the decoupling temperature $T_{d}$ of the GUT axion (for $f_{a}=10^{12}$ GeV). For $T_{\rm RH}>T_{d}$, the GUT axion is thermalized with the SM bath and thus $\Delta N_{\rm eff}$ no longer rises with $T_{\rm RH}$.
Figure 7: A comparison of our dimension-5 and dimension-6 results for the Frogatt-Neilsen textured case and for $f_{a}=10^{12}$ GeV as a function of $T_{\rm RH}$ and $\mathcal{N}$. The existing and future CMB constraints are shown as in Fig. 6. The $\Delta N_{\rm eff}$ expected are shown with (dashed) and without (solid) $D=6$ operators on for $\mathcal{N}=5$, $\mathcal{N}=10$, and $\mathcal{N}=30$. The kinks on the solid lines indicate the decoupling temperature $T_{d}\sim 3\times 10^{10}$ GeV of the dimension-6 operator (for $f_{a}=10^{12}$ GeV). For $T_{\rm RH}>T_{d}$, all the axions are thermalized and $\Delta N_{\rm eff}$ is saturated.

### v2

@@ 3.2 Couplings to Fermions
In this section, we consider an EFT where axions couple to fermions at tree-level, in addition to the gauge sector, motivated by the DFSZ scenario Dine et al. (1981); Zhitnitsky (1980). Those couplings can be written as in (18), which we reproduce here for convenience,
@@ 3.2.1 Anarchy
3.2.2 Froggatt-Nielsen Texture
We next assume that the axion couplings to the fermion sector are such that the SM flavor puzzle is solved by the Froggatt-Nielsen (FN) mechanism Froggatt and Nielsen (1979) at a scale $\Lambda_{\rm FN}\gg f_{a}$. In the FN setup, the SM is extended with a global $U(1)_{\cal H}$ horizontal symmetry which is spontaneously broken by the VEV $\langle S\rangle$ of a new complex scalar $S$. The SM Yukawas arise solely from powers of $\epsilon\equiv\langle S\rangle/\Lambda_{\rm FN}$.
This idea can be generalized by considering a texture Leurer et al. (1993, 1994), the set of charges of the SM fermions under $U(1)_{\cal H}$. In this work, we consider such an example of a texture, but we take a bottom-up approach and do not specify the UV degrees of freedom that provide such charges. We adopt a convention where $S$ has charge 1 and the SM Higgs has charge 0 under $U(1)_{\cal H}$, we can express the SM Yukawa couplings
@@ 3.2.2 Froggatt-Nielsen Texture
where $\kappa$’s are some ${\cal O}(1)$ numbers. In most implementations, the ratio $\langle S\rangle/\Lambda_{\rm FN}$ is identified with $\epsilon\equiv\sin\theta_{c}\approx 0.23$, and a texture is identified that reproduces the SM Yukawas at low energies. To conserve charge, we see that $n_{nm}^{d}=|{\cal H}(Q_{n})-{\cal H}(d_{m})|$ and $n_{nm}^{u}=|{\cal H}(Q_{n})-{\cal H}(u_{m})|$. Note that if the argument of the absolute value is negative, then instead $S^{\dagger}$ is coupled to the SM fermions. Various choices of texture are possible, see, e.g.Leurer et al. (1993, 1994). Once the FN charges are assigned to the SM fermions, they also dictate the sizes of the couplings of those fermions to other fields, including new physics scenarios like the axion couplings. In particular, the structure of the $\phi_{i}$ coupling to the SM fermions requires $c_{F,nm}\propto\epsilon^{|{\cal H}(\Psi_{F}^{n})-{\cal H}(\Psi_{F}^{m})|}$. Using the horizontal charge assignments given in Table 1 of Asadi et al. (2023), we find
@@
The numbers $X,Y$ take values $\pm 1$ and together with the other sign choices in $c_{u},c_{d},c_{e}$, there are $2^{5}$ possibilities for the FN charges. For an example benchmark, we choose X=Y=1X=-Y=-1, the positive sign in $c_{u},c_{d}$, and the negative sign in $c_{e}$; this choice is motivated by holomorphy of a SUSY UV completion of the axiverse EFT Asadi et al. (2023). This leads to
@@
Finally, we note that the angular mode of $S$ is a light pseudoscalar which could also freeze-in. In this work, we do not consider its contribution to $\Delta N_{\rm eff}$. In some models it is identified with the axion Ema et al. (2017); Calibbi et al. (2017).
@@ 3.2.3 Minimal Flavor Violation
Our final assumption for the axion-fermion couplings does introduce correlations between the axions coupled to different SM fermions, so we will now find that there are less than 42 axion-fermion couplings at dimension five. We investigate in this context the principle of Minimal Flavor Violation (MFV) D’Ambrosio et al. (2002), where the SM Yukawas transform as spurions under the $SU(3)^{5}$ flavor symmetry group. Then the axion-fermion couplings can be expanded in powers of the spurions, where $c_{F}$ is required to transform as $\mathbf{8}_{F}$ under $SU(3)_{\psi}\subset SU(3)^{5}$. It is useful to define the Hermitian matrices $X_{U}\equiv Y_{u}Y_{u}^{\dagger}\sim\mathbf{8}_{Q}$, $X_{D}\equiv Y_{d}Y_{d}^{\dagger}\sim\mathbf{8}_{Q}$, $U_{k}\equiv Y_{u}^{\dagger}X_{D}^{k-1}Y_{u}\sim\mathbf{8}_{u}$, $D_{k}\equiv Y_{d}^{\dagger}X_{U}^{k-1}Y_{d}\sim\mathbf{8}_{d}$, $L\equiv Y_{e}Y_{e}^{\dagger}\sim\mathbf{8}_{L}$, $E\equiv Y_{e}^{\dagger}Y_{e}\sim\mathbf{8}_{e}$. Using the Cayley-Hamilton theorem, we expand the couplings in only these matrices and remove related terms. We find
@@
where all expansion coefficients are expected to be $\mathcal{O}(1)$. In the limit of zero neutrino masses, the expansion in the lepton sector is complete. While we have not explicitly written the full expansion in the quark sector, it can be systematically computed Colangelo et al. (2009). In our numerical computation, we use the full expansion without the truncation shown above. It is useful to work in a basis where one of the quark Yukawa couplings is diagonal. We label the up-basis the one where $Y_{u}={\rm diag}(y_{u},y_{c},y_{t})$ and $Y_{d}=V_{\rm CKM}{\rm diag}(y_{d},y_{s},y_{b})$. The down-basis is related by a $SU(3)_{Q}$ rotation by $V_{\rm CKM}^{\dagger}$. Explicitly, in either basis $c_{u}$, $c_{d}$, $c_{L}$, and $c_{e}$ are identical, while $c_{Q}^{\rm up}=V_{\rm CKM}c_{Q}^{\rm down}V_{\rm CKM}^{\dagger}$. We choose to work in the up-basis. We have explicitly checked that working in the down-basis leads to identical results.
@@ 4.1 Effects of Mass Mixing
For simplicity, consider a two-axion system, where the axions have masses $m_{1}$ and $m_{2}$. The mass basis does not coincide with the flavor basis generically, and therefore, the axion of one flavor can oscillate into another flavor. In analogy with neutrino oscillations in vacuum, the probability of a flavor oscillation in time $t$ is given by $\sin^{2}(\vartheta)\times\sin^{2}((m_{2}^{2}-m_{1}^{2})t/(4k))$ Raffelt (1996), where $\vartheta$ is the rotation angle between the mass basis and the flavor basis, and $k$ is the 3-momenta of both the axions. Here, we have treated both axions to be relativistic. Since we are interested in the production of these axions from the thermal bath, $k\sim T$, the temperature of the SM bath. Taking $t\sim 1/H$, at $T\gg{\rm GeV}$ and $m_{2},m_{1}\ll\rm{eV}$ (as we will justify shortly),
@@ 4.2 QCD Axion Cosmology
Depending on its mass, the QCD axion can either contribute as dark matter or dark radiation. Since our EFT treatment is valid for $T_{\rm RH}<f_{a}$, we only focus on DM production through the misalignment mechanism, which implies Navas and others (2024)
@@
for misalignment angle $\theta_{i}\lesssim 1$. The QCD axion can also be produced from the SM bath, either through pion scattering or through quarks and gluons, depending on the temperature. Ref. Bianchini et al. (2024) places an upper bound $m_{\rm QCD}\leq 0.16$ eV at 95% CL, implying $f_{a}\geq 3.6\times 10^{7}$ GeV, by computing $a_{\rm QCD}$ production from pion scattering at temperatures $\lesssim 150$ MeV, the QCD crossover temperature (see also Ref. D’Eramo et al. (2022a) for an analysis with higher initial temperatures, with similar results). Above this temperature, but below the electroweak scale, perturbative computations are not fully trustworthy. See Refs. Grilli di Cortona et al. (2016); D’Eramo et al. (2022b, 2021); Notari et al. (2023) for a discussion of the regime of validity of perturbation theory of QCD at finite temperature. However, convergence of perturbative QCD improves at higher temperatures, and we expect the associated rates to be under better control for $T_{\rm RH}\gtrsim$ TeV, which is where we focus on in this work.
@@ 4.4.1 Production via Gauge Couplings
To estimate axion production from gauge couplings, we use the production rates based on the thermal field theory calculations in Salvio et al. (2014) (for calculations at zero temperature, see Masso et al. (2002)). These rates also determine the axion decoupling temperatures $T_{d}$, which is when the axion production rate equals the Hubble expansion rate. We evaluate $T_{d}$ for the three gauge interactions and show the result in Fig. 1, including the GUT scenario. We also use these thermal rates to obtain an approximate abundance of $\Delta N_{\rm eff}$,
@@ 4.4.2 Production via Fermionic Couplings
The associated freeze-in production can be computed analytically via solving the full Boltzmann equation for the axion phase space distribution function. [^3] For a general numerical treatment see, e.g.D’Eramo et al. (2024); Badziak and Laletin (2025); D’Eramo and Lenoci (2024). We can express the relevant processes for freeze-in production as,
@@
where we have defined a momentum-dependent ‘production rate’ $\Gamma(p,t)$. To solve this we note the physical momentum $p=q/a$ where $q$ is the comoving momentum. We then express $\mathfrak{f}_{a}(p,t)=\tilde{\mathfrak{f}}_{a}(q,a)$ and analogously for the other quantities. The Boltzmann equation then becomes a collection of ordinary differential equations for each comoving momentum D’Agnolo et al. (2017),
@@ 4.4.3 Production via Dim-6 Higgs Coupling
However, the fact that there are two axions in the final state leads to a nonlinear integro-differential equation which must be solved numerically. We leave this to future work, and make the approximation that we can ignore the backwards process: $\mathfrak{f}_{a}(\boldsymbol{k}_{1},t)\mathfrak{f}_{a}(\boldsymbol{k}_{2},t)=0$. Clearly, this can only be justified in the regime that the axion does not thermalize; we ceiling $\Delta N_{\rm eff}$ at a thermal abundance. Following Edsjo and Gondolo (1997); Hall et al. (2010); Blennow et al. (2014), one can then solve the (integrated) Boltzmann equations for the axion number density,
@@
While this result matches parametrically with Ref. Bauer et al. (2023), its numerical value differs by several orders of magnitude. For a verification, we have checked that our computation matches with micrOMEGAs Alguero et al. (2024) numerical output.
@@ 5 Results for Benchmark EFTs
We also show the sensitivity of current and upcoming CMB experiments. The horizontal lines correspond to existing one-sided 95% upper limits $\Delta N_{\rm eff}^{95\%}$. SPT-Planck-ACT (SPA) corresponds to the constraint from the recent SPT-3G analysis Camphuis and others (2025) when floating the helium abundance $Y_{P}$ in the fit, yielding $N_{\rm eff}=2.99_{-0.26}^{+0.22}$, so that $\Delta N_{\rm eff}^{95\%}\approx 0.31$. We note that without floating $Y_{P}$, we find $\Delta N_{\rm eff}^{95\%}\approx-0.04$ due to the low central values from Planck and ACT. For this reason, we adopt the more conservative result from floating $Y_{P}$. Indeed, the ACT results Calabrese and others (2025) (including Planck lensing and BOSS BAO) yield $\Delta N_{\rm eff}^{95\%}\approx 0.04$. We show instead the result from restricting to physical parameter space with $\Delta N_{\rm eff}>0$: $\Delta N_{\rm eff}^{95\%}=0.17$, but in plots we shade only the region excluded by the SPA analysis. Finally, we show projections for Simons Observatory (SO), CMB-S4 (S4), and CMB-HD (HD) Abitbol and others (2025); Abazajian and others (2016); MacInnis et al. (2024).
@@
Figure 6: A summary of our results considering only dimension-5 couplings for $f_{a}=10^{12}$ GeV as a function of $T_{\rm RH}$ and different assumptions about the axion couplings. The existing and future CMB constraints are shown as horizontal lines: the combination of SPT-3G, Planck, and ACT (SPA, black solid), ACT alone (gray solid), SO (dashed), S4 (dot-dashed), and HD (dotted). The $\Delta N_{\rm eff}$ expected are also shown assuming anarchic couplings, a Froggatt-Nielsen texture, MFV, hadronic, and a hadronic GUT. The kink on the hadronic GUT (blue) line around $2\times 10^{11}$ GeV indicates the decoupling temperature $T_{d}$ of the GUT axion (for $f_{a}=10^{12}$ GeV). For $T_{\rm RH}>T_{d}$, the GUT axion is thermalized with the SM bath and thus $\Delta N_{\rm eff}$ no longer rises with $T_{\rm RH}$.
Figure 7: A comparison of our dimension-5 and dimension-6 results for the Froggatt-Nielsen textured case and for $f_{a}=10^{12}$ GeV as a function of $T_{\rm RH}$ and $\mathcal{N}$. The existing and future CMB constraints are shown as in Fig. 6. The $\Delta N_{\rm eff}$ expected are shown with (dashed) and without (solid) $D=6$ operators on for $\mathcal{N}=5$, $\mathcal{N}=10$, and $\mathcal{N}=30$. The kinks on the solid lines indicate the decoupling temperature $T_{d}\sim 3\times 10^{10}$ GeV of the dimension-6 operator (for $f_{a}=10^{12}$ GeV). For $T_{\rm RH}>T_{d}$, all the axions are thermalized and $\Delta N_{\rm eff}$ is saturated.